(11 points) Graduate School. Jaime is applying for graduate school at a local university. In the past, 43% of the applicants to this university have been accepted. It is also known that 69% of those students who have been accepted have had GMAT scores in excess of 550 while 38% of the students who were not accepted had GMAT scores in excess of 550. Use this information and a theoretical group of 10,000 applicants to construct a table for this admissions scenario. Round all answers and results to four decimal places. GMAT> 550 GMAT 550 Total Accepted 2967 Not Accepted 2166 Total 5133 4300 1333 3534 4867 5700 10000 a. What percentage of applicants to this university score over 550 on the GMAT? 51.33 b. What percentage of applicants who score over 550 on the GMAT would we expect to be accepted to this university?29.67 c. What percentage of applicants who do not score over 550 on the GMAT would we expect to be accepted to this university? 13.33
Answer #1

from given data:

GMAT>550 GMAT<=550 total
accepted 2967 1333 4300
not accepted 2166 3534 5700
total 5133 4867 10000

a)% over 550 on the GMAT =(5133/10000)*100 =51.3300%

b) P(accepted |over 550)=(2967/5133)*100=57.8025 %

c) P(accepted |<=550)=(1333/4867)*100=27.3885%